# Discriminative Adversarial Unlearning

### 10 Feb 2024 v1

@@ 4.1 Criteria
  • The general capacity of the unlearned model to defend against MIA. For this metric we employ a separate prediction confidence based attack, which utilizes the differences in confidence of the trained network for samples arising from $\mathcal{D}$ vs $\mathcal{D}_{t}$ or $\mathcal{D}_{v}$. The confidence based predictor [59, 52] is formulated through training a support vector machine on the prediction confidences of the network for samples arising from $\mathcal{D}$ and $\mathcal{D}_{t}$ and subsequently used to predict the origin of $\mathcal{D}_{f}$. A successful attack will therefore perform with high MIA-Efficacy on $\mathcal{D}_{f}$. Formally, for an attacker to fail, it must predict samples arising from $\mathcal{D}_{f}$ as true negatives and therefore, $\mathbf{MIA\text{-}efficacy}=TN/|\mathcal{D}_{f}|$ where $TN$ denotes the true negative inferences by the attacker.
@@
  • Fine-tuning (FT) [58, 23]: This method involves finetuning the network only on the retain set $\mathcal{D}_{r}$ in order to induce amnesia over the network’s prior training on $\mathcal{D}_{f}$ leveraging the network’s tendency for catastrophic forgetting, and ensconce the performance over the retain set.
  • Gradient-ascent (GA) [25, 54]: Using the negative of the classification loss evaluated on $\mathcal{D}_{f}$, this method retrains the model in order to explicitly negate the prior training and induce forgetting.
  • Fisher Forget (FF) [4, 32]: FF uses estimates the hessian of the defender network using fisher information matrix, in order to informatively corrupt the of the pre-trained network’s parameters $\theta_{0}$ using additive gaussian noise in the image space.
  • Influence Unlearning (IU) [35]: This method uses estimates of influence functions which evaluate the change in parameters of a pre-trained network due to the presence/absence of samples in $\mathcal{D}$ through ”upweighing” [12] these samples on the network parameters, allowing for the method to manipulate the influence of the loss due to $\mathcal{D}_{f}$ on the defender.
@@ 4.2 Configuration
Training is conducted on a single Nvidia RTX-A6000 GPU. The defender’s base model is the ResNet-18 [27], while the attacker network adopts a transformer structure with $3$ multihead self-attention layers and $4$ attention heads, retaining the default configuration from Yuan and Zhang [61]. Pre-training of the defender employs a learning rate of $0.1$ for both Cifar datasets. For random forgetting, the (un-)learning rate is set to $0.01$ with the dense model, while a higher rate of $0.03$ is found to be more suitable for a sparse model in this scenario. In the class-wise unlearning scenario, a rate of $0.02$ is used for both levels of sparsity and datasets. Empirical adjustments are made to the $\alpha$ and $\beta$ parameters, setting them to $0.9$ and $0.001$, respectively.
@@ 4.3 Results
Our results are listed in the Table 1 for all the aforementioned configurations. We conduct our evaluations under the random forgetting scheme where $10\%$ of the samples are chosen randomly to be scrubbed from the pre-trained model, and class-wise forgetting scheme where all samples from a randomly chosen class are tasked for removal. All methods were evaluated using the dense model and a $95\%$ sparse model, in order to gauge the robustness and efficacy of the listed methods under varying conditions of sparsity. Sparsity in the networks is enforced using One-Shot Magnitude Pruning (OMP) [41] due to its lower computational overhead and better generalization [33]. The gold-standard for unlearning which is to retrain the model on $\mathcal{D}_{r}$ is also evaluated. The assessment employs the metrics outlined in section 4.1. It is imperative to emphasize that elevated values across these metrics do not inherently denote superior performance and unlearning methods should exhibit a minimal disparity in comparison to the retrain method. Consequently, our evaluation incorporates the computation of the average disparity of scores, aiming to simulate a measure of proximity with this gold standard. The experiments reveal that our proposed method surpasses all baseline approaches across the utilized metrics and attains the utmost proximity to the gold standard. In the context of the random forgetting scheme, our method consistently achieves the closest performance to the retrain method. Simultaneously, it maintains superior performance in both the retain set (RA) and test set (TA). Contrastingly, other baseline methods exhibit ineffectiveness in the task of forgetting, resulting in excessively high scores for RA and TA, or they adopt over-compensatory measures, causing a decline in the model’s overall classification capacity,, as evidenced by the MIA-efficacy metric. Notably, FF achieves significant scores in Cifar-10 under this scheme but experiences a considerable decline with over $5\%$ point gap from the retrain benchmark in the TA and RA metrics.
@@
Sparse models however, ease the process of unlearning performance for the methods as noted by Jia et al. [33] whilst incurring a marginal cost over the TA and RA metrics. This phenomenon is particularly pronounced in class-wise forgetting schemes, where the baseline methods demonstrate significant improvements across the metrics and resulting in substantially lower Avg. Disparity scores. Our approach demonstrates a comparable trend, closely aligning with the retraining method across various metrics. Notably, we argue that incorporating sparsity as a performance enhancement measure yields marginal benefits in our context. The algorithm consistently achieves robust performance with dense variants, presenting challenges for the baselines, even when employing $95\%$ sparse models. Moreover, $\mathcal{V}_{ss}$ plays a critical role in enhancing performance as we elucidate in Section 4.5.
@@ 4.4 MIA-Robustness
  • Prediction correctness MIA [59]: The attacker is trained to infer membership of a record if the defender’s predictions are correct. This attack capitalizes on the defender’s ability to predict accurately on $\mathcal{D}$ which may not generalize to $\mathcal{D}_{t}$.
  • Prediction confidence MIA [48]: The attacker infers membership based on the maximum prediction confidence scores of the defender across all samples. This leverages differential scores of confidence measures for the defender over $\mathcal{D}$ and $\mathcal{D}_{t}$.
  • Prediction entropy MIA [48]: The attacker evaluates the entropy of predictions over samples from $\mathcal{D}$ and $\mathcal{D}_{t}$, considering the possibility that the entropy of predictions for $\mathcal{D}_{t}$ is higher.
  • Modified prediction entropy MIA [52]: In addition to prediction entropy, the attacker incorporates ground truth label information to leverage the defender’s capacity for false negatives, inferring membership based on, for example, incorrectly predicted records.
@@ 5 Conclusion
In this study, we employ a straightforward and differentiable framework, adhering to the well-established principles of min-max optimization, to address the challenge of machine unlearning. This framework demonstrates high efficiency, yielding substantial runtime improvements over retraining (at least 5x in the case of Cifar-100 class-wise). Moreover, its effectiveness is evident through consistently achieving the lowest disparity scores across various configurations. Remarkably, the proposed method attains near-perfect performance, particularly notable in the class-wise forgetting schemes. In these scenarios, the unlearning accuracy and MIA robustness reach 100% for Cifar-10 and approach this level for Cifar-100, even within dense network regimes with minimal compromises observed in the test accuracies. Furthermore, the method avoids the expenses of hessian matrix approximations for deep networks and involves minimal parameter tuning requirements. The approach additionally, is adaptable to incorporation of advanced MIA attack techniques and improvements over the objectives of the defender and attacker in order to ensure better performances and convergence. Nevertheless, it is noteworthy that the framework excels under the class-wise unlearning scheme, leaving room for refinement in the context of the random unlearning scheme—potentially a more realistic scenario for machine unlearning. We hypothesize that integrating improved optimization procedures [2, 26, 34] and employing stronger, multiple attackers leveraging additional information from predictions [59, 48, 51] could enhance the robustness of the framework in this scheme. This avenue remains a subject for future exploration.

### 13 Feb 2024 v2

@@ 4.1 Criteria
  • The general capacity of the unlearned model to defend against MIA. For this metric we employ a separate prediction confidence based attack, which utilizes the differences in confidence of the trained network for samples arising from $\mathcal{D}$ vs $\mathcal{D}_{t}$ or $\mathcal{D}_{v}$. The confidence based predictor (Yeom et al., 2018; Song et al., 2019) is formulated through training a support vector machine on the prediction confidences of the network for samples arising from $\mathcal{D}$ and $\mathcal{D}_{t}$ and subsequently used to predict the origin of $\mathcal{D}_{f}$. A successful attack will therefore perform with high MIA-Efficacy on $\mathcal{D}_{f}$. Formally, for an attacker to fail, it must predict samples arising from $\mathcal{D}_{f}$ as true negatives and therefore, $\mathbf{MIA\text{-}efficacy}=TN/|\mathcal{D}_{f}|$ where $TN$ denotes the true negative inferences by the attacker.
@@
  • Fine-tuning (FT) (Warnecke et al., 2021; Golatkar et al., 2020): This method involves finetuning the network only on the retain set $\mathcal{D}_{r}$ in order to induce amnesia over the network’s prior training on $\mathcal{D}_{f}$ leveraging the network’s tendency for catastrophic forgetting, and ensconce the performance over the retain set.
  • Gradient-ascent (GA) (Graves et al., 2021; Thudi et al., 2022): Using the negative of the classification loss evaluated on $\mathcal{D}_{f}$, this method retrains the model in order to explicitly negate the prior training and induce forgetting.
  • Fisher Forget (FF) (Becker and Liebig, 2022; Izzo et al., 2021): FF uses estimates the hessian of the defender network using fisher information matrix, in order to informatively corrupt the of the pre-trained network’s parameters $\theta_{0}$ using additive gaussian noise in the image space.
  • Influence Unlearning (IU) (Koh and Liang, 2017): This method uses estimates of influence functions which evaluate the change in parameters of a pre-trained network due to the presence/absence of samples in $\mathcal{D}$ through ”upweighing” (Cook and Weisberg, 1980) these samples on the network parameters, allowing for the method to manipulate the influence of the loss due to $\mathcal{D}_{f}$ on the defender.
@@ 4.2 Configuration
Training is conducted on a single Nvidia RTX-A6000 GPU. The defender’s base model is the ResNet-18 (He et al., 2016), while the attacker network adopts a transformer structure with $3$ multihead self-attention layers and $4$ attention heads, retaining the default configuration from Yuan and Zhang (2022). Pre-training of the defender employs a learning rate of $0.1$ for both Cifar datasets. For random forgetting, the (un-)learning rate is set to $0.01$ with the dense model, while a higher rate of $0.03$ is found to be more suitable for a sparse model in this scenario. In the class-wise unlearning scenario, a rate of $0.02$ is used for both levels of sparsity and datasets. Empirical adjustments are made to the $\alpha$ and $\beta$ parameters, setting them to $0.9$ and $0.001$, respectively.
@@ 4.3 Results
Our results are listed in the Table 1 for all the aforementioned configurations. We conduct our evaluations under the random forgetting scheme where $10\%$ of the samples are chosen randomly to be scrubbed from the pre-trained model, and class-wise forgetting scheme where all samples from a randomly chosen class are tasked for removal. All methods were evaluated using the dense model and a $95\%$ sparse model, in order to gauge the robustness and efficacy of the listed methods under varying conditions of sparsity. Sparsity in the networks is enforced using One-Shot Magnitude Pruning (OMP) (Ma et al., 2021) due to its lower computational overhead and better generalization (Jia et al., 2023). The gold-standard for unlearning which is to retrain the model on $\mathcal{D}_{r}$ is also evaluated. The assessment employs the metrics outlined in section 4.1. It is imperative to emphasize that elevated values across these metrics do not inherently denote superior performance and unlearning methods should exhibit a minimal disparity in comparison to the retrain method. Consequently, our evaluation incorporates the computation of the average disparity of scores, aiming to simulate a measure of proximity with this gold standard. The experiments reveal that our proposed method surpasses all baseline approaches across the utilized metrics and attains the utmost proximity to the gold standard. In the context of the random forgetting scheme, our method consistently achieves the closest performance to the retrain method. Simultaneously, it maintains superior performance in both the retain set (RA) and test set (TA). Contrastingly, other baseline methods exhibit ineffectiveness in the task of forgetting, resulting in excessively high scores for RA and TA, or they adopt over-compensatory measures, causing a decline in the model’s overall classification capacity,, as evidenced by the MIA-efficacy metric. Notably, FF achieves significant scores in Cifar-10 under this scheme but experiences a considerable decline with over $5\%$ point gap from the retrain benchmark in the TA and RA metrics.
@@
Sparse models however, ease the process of unlearning performance for the methods as noted by Jia et al. (2023) whilst incurring a marginal cost over the TA and RA metrics. This phenomenon is particularly pronounced in class-wise forgetting schemes, where the baseline methods demonstrate significant improvements across the metrics and resulting in substantially lower Avg. Disparity scores. Our approach demonstrates a comparable trend, closely aligning with the retraining method across various metrics. Notably, we argue that incorporating sparsity as a performance enhancement measure yields marginal benefits in our context. The algorithm consistently achieves robust performance with dense variants, presenting challenges for the baselines, even when employing $95\%$ sparse models. Moreover, $\mathcal{V}_{ss}$ plays a critical role in enhancing performance as we elucidate in Section 4.5.
@@ 4.4 MIA-Robustness
  • Prediction correctness MIA (Yeom et al., 2018): The attacker is trained to infer membership of a record if the defender’s predictions are correct. This attack capitalizes on the defender’s ability to predict accurately on $\mathcal{D}$ which may not generalize to $\mathcal{D}_{t}$.
  • Prediction confidence MIA (Salem et al., 2018): The attacker infers membership based on the maximum prediction confidence scores of the defender across all samples. This leverages differential scores of confidence measures for the defender over $\mathcal{D}$ and $\mathcal{D}_{t}$.
  • Prediction entropy MIA (Salem et al., 2018): The attacker evaluates the entropy of predictions over samples from $\mathcal{D}$ and $\mathcal{D}_{t}$, considering the possibility that the entropy of predictions for $\mathcal{D}_{t}$ is higher.
  • Modified prediction entropy MIA (Song et al., 2019): In addition to prediction entropy, the attacker incorporates ground truth label information to leverage the defender’s capacity for false negatives, inferring membership based on, for example, incorrectly predicted records.
@@ 5 Conclusion
In this study, we employ a straightforward and differentiable framework, adhering to the well-established principles of min-max optimization, to address the challenge of machine unlearning. This framework demonstrates high efficiency, yielding substantial runtime improvements over retraining (at least 5x in the case of Cifar-100 class-wise). Moreover, its effectiveness is evident through consistently achieving the lowest disparity scores across various configurations. Remarkably, the proposed method attains near-perfect performance, particularly notable in the class-wise forgetting schemes. In these scenarios, the unlearning accuracy and MIA robustness reach 100% for Cifar-10 and approach this level for Cifar-100, even within dense network regimes with minimal compromises observed in the test accuracies. Furthermore, the method avoids the expenses of hessian matrix approximations for deep networks and involves minimal parameter tuning requirements. The approach additionally, is adaptable to incorporation of advanced MIA attack techniques and improvements over the objectives of the defender and attacker in order to ensure better performances and convergence. Nevertheless, it is noteworthy that the framework excels under the class-wise unlearning scheme, leaving room for refinement in the context of the random unlearning scheme—potentially a more realistic scenario for machine unlearning. We hypothesize that integrating improved optimization procedures (Arjovsky et al., 2017; Gulrajani et al., 2017; Kodali et al., 2017) and employing stronger, multiple attackers leveraging additional information from predictions (Yeom et al., 2018; Salem et al., 2018; Song and Mittal, 2021) could enhance the robustness of the framework in this scheme. This avenue remains a subject for future exploration.